Air is pumped into a spherical balloon such that the surface area of the balloon increases at a rate of 10 cm2 per second. Find the rate at which the radius is increasing when the radius is 4 cm. [ i.e. find when r = 4] [Note : surface area, A = 4πr2 ] 


1
Expert's answer
2021-09-09T00:49:27-0400

Let the rate of area change be "\dfrac{dA}{dt} = 10cm^2/s". Assuming that "A = 4\pi r^2" and applying chain rule, obtain:


"\dfrac{d}{dt}(4\pi r^2) = 4\pi\cdot 2r\cdot \dfrac{dr}{dt} = 10cm^2/s"

where "\dfrac{dr}{dt}" is the rate at which the radius is increasing. Expressing this rate and substituing "r = 4cm" (current radius), obtain:


"\dfrac{dr}{dt} = \dfrac{10cm^2/s}{8\pi r} = \dfrac{10cm^2/s}{8\pi \cdot 4cm} \approx 0.1cm/s"

Answer. 0.1 cm/s.